Can you create one formula for this question instead of 2?
if 10 jelly beans weigh 12 grams, how many jelly beans in 650 grams

Answer by Theo(13342)

x = 650 * 10 / 12

x represents the number of jelly beans that weigh 650 grams

that equation comes from the ratio:

10 / 12 = x / 650

multiply both sides of this equation by 650 and you get:

650 * 10 / 12 = x which is the same as:

x = 650 * 10 / 12.

the answer is 541 and 2/3 jelly beans will weigh 650 grams.

divide 650 by 541 and 2/3 and you will get 1.2 graphs per jelly bean.

this is correct since 10 jelly beans weigh 12 grams which also comes out to be 1.2 grams per jelly bean.

the equation is a proportion equationm that was then used to solve for x.

the proportion was 10 / 12 = x / 650

12 / 10 = 650 / x would have worked just as well.

either equation gets you x = 541 and 2/3 jelly beans.